An ideal gas at initial temperature (T) and initial volume(V) is expanded adiabatically to a volume (2V). The gas is then expanded isothermally a volume (5V) and thereafter compressed adiabatically so that the temperature of the gas becomes again (T).If the final volume of the gas is x(V). then the value of x is......
1
Expert's answer
2020-01-21T07:39:48-0500

The combined gas law:

"\frac{p \cdot V}{T} = const",

where "p" is the pressure of the gas, "V" is the volume of the gas, "T" is the temperature of the gas.

Initial parameters: "p_1=P", "V_1=V" and "T_1=T".

1) adiabatic process is characterized by the constant entropy.

Gas can be characterized by the parameters "p_2", "V_2" and "T_2".

"V_2=2V"

"T_2=(\frac{V_2}{V_1})^{1 - \gamma} \cdot T_1 = (\frac{2V}{V})^{1 - \gamma} \cdot T= (2)^{1 - \gamma} \cdot T"

where "\gamma" is the heat capacity ratio, which is constant for a calorifically perfect gas.

2) isothermal process is characterized by the constant temperature.

Gas can be characterized by the parameters "p_3", "V_3" and "T_3".

"V_3=5V"

"T_3=T_2= (2)^{1 - \gamma} \cdot T"

3) adiabatic process is characterized by the constant entropy.

Gas can be characterized by the parameters "p_4", "V_4" and "T_4".

"T_4=T", so "\frac{T_4}{T_3}=\frac{T}{(2)^{1 - \gamma} \cdot T} = (2)^{ \gamma - 1}"

"V_4=V_3 (\frac{T_4}{T_3})^{\frac{1}{1- \gamma}} = 5V ((2)^{ \gamma - 1})^{\frac{1}{1- \gamma}}=(2)^{\frac{\gamma -1}{1 - \gamma}} \cdot 5V= (2)^{-1} \cdot 5V = \frac{5V}{2} = 2.5V"


Answer: If the final volume of the gas is "x(V)", then the value of "x" is 2.5.